OpenStudy (anonymous):

Evaluate the expression for the given value of the variable(s). -x2 - 4x; x = –3 B. –17 C. 11 D. 3

whitemonsterbunny17 (whitemonsterbunny17):

Just replace "x" with -3

whitemonsterbunny17 (whitemonsterbunny17):

When you replace "x" with -3, what does your problem look like?

whitemonsterbunny17 (whitemonsterbunny17):

@bvbarmy17

OpenStudy (anonymous):

-(3)2-4(-3);x=-3

OpenStudy (anonymous):

@Whitemonsterbunny17

whitemonsterbunny17 (whitemonsterbunny17):

Yes. So you know PEMDAS?

OpenStudy (anonymous):

yes

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so tell me what PEMDAS stands for.

OpenStudy (anonymous):

Parentheses Exponents Multplicaton Divide Add Subtract

whitemonsterbunny17 (whitemonsterbunny17):

Basically, yes. \(\normalsize\bf\color{red}P\color{gold}E\color{lime}M\color{cyan}D\color{blue}A\color{purple}S\): \(\normalsize\bf\color{red}P\)arentheses \(\normalsize\bf\color{gold}E\)xponents \(\normalsize\bf\color{lime}M\)ultiplication \(\normalsize\bf\color{cyan}D\)ivision \(\normalsize\bf\color{blue}A\)ddition \(\normalsize\bf\color{purple}S\)ubtraction We have no real parentheses, so we move on. There are also no exponents, so next is: PE\(\normalsize\bf\color{lime}M\)DAS \(\normalsize\bf\color{lime}M\)ultiplication

OpenStudy (anonymous):

ok

whitemonsterbunny17 (whitemonsterbunny17):

So go ahead, and do the \(\normalsize\bf\color{lime}M\)ultiplication. (:

OpenStudy (anonymous):

-6-12

OpenStudy (anonymous):

-6-12=-18

whitemonsterbunny17 (whitemonsterbunny17):

Yeah. I got the same thing.

OpenStudy (anonymous):

but its not answer c hoice

whitemonsterbunny17 (whitemonsterbunny17):

Maybe they had a typo and meant -18.

whitemonsterbunny17 (whitemonsterbunny17):

Instead of -17

OpenStudy (anonymous):

yea maybe

OpenStudy (anonymous):

ill pick "B" its the closes

whitemonsterbunny17 (whitemonsterbunny17):

Okay, tell me if it was right. (:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

they said it was wrong

whitemonsterbunny17 (whitemonsterbunny17):

Really? >.<

OpenStudy (anonymous):

yea idk what it is then

whitemonsterbunny17 (whitemonsterbunny17):

No matter what I try, I'm still getting -18 as the result.

OpenStudy (anonymous):

-(-3)2-4(-3)= 6+12= +18